# Stable Neural Stochastic Differential Equations in Analyzing Irregular Time Series Data

### v2

@@ 2 Related Work & Preliminaries
Let $(\Omega,{\mathcal{F}},P)$ be a probability space. Fix integer $d\geq 1$. For an ${\mathbb{R}}^{d}$-valued random variable $X$, its law is denoted by ${\mathcal{L}}(X)$. $|\cdot|$ is the Euclidean norm where the dimension of the space may vary depending on the context. We denote by $L^{p}(\Omega)$ the set of $X$ with $(\mathbb{E}[|X|^{p}])^{1/p}<\infty$. Let ${\mathcal{P}}({\mathbb{R}}^{d})$ denote the set of probability measures on ${\mathcal{B}}({\mathbb{R}}^{d})$. For two probability measures $\mu,\nu\in{\mathcal{P}}({\mathbb{R}}^{d})$, let ${\mathcal{C}}(\mu,\nu)$ denote the set of probability measures $\Pi$ on ${\mathcal{B}}({\mathbb{R}}^{2d})$ such that its respective marginals are $\mu$ and $\nu$. Then, the Wasserstein distance of order $p\geq 1$ is defined as
@@ 3.1 The proposed Neural SDEs
where the initial condition $h(\cdot;\theta_{h})$ is an affine function with parameter $\theta_{h}$, the drift function $\gamma(\cdot,\cdot;\theta_{\gamma})$ is a neural network with parameter $\theta_{\gamma}$ and the diffusion function $\sigma(\cdot;\theta_{\sigma})$ is a neural network with parameter $\theta_{\sigma}$. Due to its exponential form, Neural GSDE has distinct characteristics associated with deep ReLU networks such as a unique positive solution and an absorbing state of 0, which are not observed in Neural LSDE and Neural LNSDE (see Proposition 3.4 (ii)).
@@ 3.2 Properties of the proposed Neural SDEs
    Assume that the activation function in $\gamma$ is either a tanh or sigmoid function. Then, the SDE in Equation (6) has a unique strong solution ${\bm{z}}(t)$, which is nonnegative almost surely, i.e., $P({\bm{z}}(t)\geq 0\mbox{ for all $t\geq 0$})=1$ a.s. Furthermore, state 0 is an absorbing state.
The proof for Proposition 3.4 can be found in Appendix A. Intuitively, Proposition 3.4 (ii) implies that the latent representation ${\bm{z}}(t)$ of the Neural GSDE has always nonnegative values. Furthermore, once the one of values of the latent representation reaches 0, it remains there forever. This property is consistent with that of ReLU networks which have positive values for activated neurons and zeros for deactivated neurons. In particular, the fact that deactivated neurons never turn to activated neurons corresponds to the property of state 0 being an absorbing state.
@@ 3.3 Robustness under distribution shift
[…]
@@
$)\displaystyle{\mathcal{W}}_{1}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}$)) $\leq\sqrt{3}L_{F}L_{h}c_{1}e^{-c_{2}T}\sqrt{\left(5+2\mathbb{E}[|{\bm{x}}|^{4}]+2\mathbb{E}[|\widetilde{\bm{x}}|^{4}]]\right)}\rho,$
$)\displaystyle{\mathcal{W}}_{2}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}$)) $\leq\sqrt{2\sqrt{3}L_{h}c_{1}}L_{F}e^{-c_{2}T/2}\left(5+2\mathbb{E}[|{\bm{x}}|^{4}]+2\mathbb{E}[|\widetilde{\bm{x}}|^{4}]]\right)^{1/4}\sqrt{\rho},$
@@ Results.
Methods Mean Squared Error ($\times 10^{-3}$)
@@ Acknowledgement
where $\gamma_{i}$ is the $i$-th element of $\gamma$ and $(\sigma(t;\theta_{\sigma}))_{ij}$ represents the $(i,j)$-th element of $\sigma(t;\theta_{\sigma})$, and ${\bm{y}}_{i}(t),{\bm{y}}_{i}(t)]$ is the quadratic variation of ${\bm{y}}_{i}(t)$ up to $t$ for $i=1,2,\ldots,d_{z}$. In the matrix form, ${\bm{z}}(t)$ satisfies
@@
Therefore, ${\bm{z}}(t)$ is the solution of Equation (6). Let $T_{0}^{i}=\inf\{t>0|z_{i}(t)=0\}$ be the first hitting time to state 0 with $z_{i}(0)>0$ for $i=1,2,\ldots,d_{z}$ where $z_{i}(t)$ is the $i$-th component of ${\bm{z}}(t)$. Since ${\mathrm{d}}z_{i}(t)=0$ for $t\geq T_{0}^{i}$, then $z_{i}(t)=0$ for all $t\geq T_{0}^{i}$. This implies that once $z_{i}(t)$ reaches 0, it remains there forever. Therefore, 0 is an absorbing state. ∎
@@
$\mbox{(Neural LSDE)}\quad{\mathrm{d}}{\bm{z}}(t)$ $=\overline{\gamma}_{1}({\bm{z}}(t);\theta_{\gamma},\theta_{\zeta}){\mathrm{d}}t+\sigma(t;\theta_{\sigma}){\mathrm{d}}W(t),$ (12)
$\mbox{(Neural LNSDE)}\quad{\mathrm{d}}{\bm{z}}(t)$ $=\overline{\gamma}_{2}(t,{\bm{z}}(t);\theta_{\gamma},\theta_{\zeta}){\mathrm{d}}t+\sigma(t;\theta_{\sigma}){\bm{z}}(t){\mathrm{d}}W(t),$ (13)
$\mbox{(Neural GSDE)}\quad\frac{{\mathrm{d}}{\bm{z}}(t)}{{\bm{z}}(t)}$ $=\overline{\gamma}_{2}(t,{\bm{z}}(t);\theta_{\gamma},\theta_{\zeta}){\mathrm{d}}t+\sigma(t;\theta_{\sigma}){\mathrm{d}}W(t).$ (14)
@@ B.1 Background on stability of SDEs
$)\displaystyle{\mathcal{W}}_{1}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}$)) $={\mathcal{W}}_{1}\left({\mathcal{L}}(F({\bm{z}}(T))),{\mathcal{L}}(F(\widetilde{\bm{z}}(T))\right)$
[…]
@@
$)\displaystyle\leq L_{h}c_{1}e^{-c_{2}T}w_{1,2}({\mathcal{L}}({\bm{x}}),{\mathcal{L}}(\widetilde{\bm{x}}$))
[…]
@@
Now focus on deriving the upper bound for $){\mathcal{W}}_{2}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}$)). Observe that
$)\displaystyle{\mathcal{W}}_{2}^{2}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}$)) $={\mathcal{W}}_{2}^{2}\left({\mathcal{L}}(F({\bm{z}}(T))),{\mathcal{L}}(F(\widetilde{\bm{z}}(T))\right)$
[…]
@@
$)\displaystyle\leq 2L_{h}c_{1}e^{-c_{2}T}w_{1,2}({\mathcal{L}}({\bm{x}}),{\mathcal{L}}(\widetilde{\bm{x}}$))
[…]
@@ B.3 Linear Noise SDE and Geometric SDE
$)\displaystyle{\mathcal{W}}_{1}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}$)) $={\mathcal{W}}_{1}\left({\mathcal{L}}(F({\bm{z}}(T))),{\mathcal{L}}(F(\widetilde{\bm{z}}(T))\right)$
[…]
@@ D.2 Network Architecture
In practice, we implemented several design choices to optimize the model’s performance and stability. Firstly, we utilized the idea of sinusoidal positional encoding, as proposed by Vaswani et al. (2017), for the time variable, ensuring that each time step possesses a unique encoding. Secondly, we opted for the tanh nonlinearity as the final operation for drift, diffusion, and all other vector fields, following the recommendation by Kidger et al. (2020). This choice is intended to prevent the model from experiencing issues related to excessively large values or gradients. Lastly, we employed layer-wise learning rates for the model’s final layer (e.g. $\times 100$), facilitating more precise and adaptive learning for the classification task. This approach can result in improved classification outcomes, as each layer can adjust its learning rate according to the complexity of the features it aims to capture.
@@ D.3 Training strategy
[…]
@@ Appendix E Detailed results of the benchmark datasets
For the proposed methodology, the training process spans $300$ epochs, employing a batch size of $64$ and a learning rate of $0.001$. To train our models on a dataset comprising irregularly sampled time series, we adopt a strategy from Shukla & Marlin (2021). This involves the modified VAE training method, where we optimize a normalized variational lower bound of the log marginal likelihood, grounded on the evidence lower bound (ELBO). Hyperparameter optimization is conducted through a grid search, focusing on the number of layers in vector field $n_{l}$ of $\left\{16,32,64,128\right\}$ and hidden vector dimensions $n_{h}$ of $\left\{16,32,64,128\right\}$. The optimal hyperparameters for the proposed methods are remarked in bold within Table 10, encompassing observed data ranging from 50% to 90%.
@@
For the proposed methodology, we spanned $500$ epochs with a batch size of $1024$ and a learning rate set at $0.001$. We performed hyperparameter optimization using a grid search, focusing specifically on the number of layers in vector field $n_{l}$ of $\left\{16,32,64,128\right\}$ and hidden vector dimensions $n_{h}$ of $\left\{16,32,64,128\right\}$. The optimal hyperparameters, as identified for our methods, are highlighted in bold in Table 11 for regular datasets. For scenarios with 30%, 50%, and 70% missing data, we applied the same hyperparameters.
@@
In the proposed method, we train for $200$ epochs with a batch size of $1024$ and a learning rate of $0.001$. The hyperparameters are optimized by grid search in the number of layers in vector field $n_{l}$ of $\left\{16,32,64,128\right\}$ and hidden vector dimensions $n_{h}$ of $\left\{16,32,64,128\right\}$. The best hyperparameters are highlighted with bold font in Table 13 and 14. Due to varying input channel counts in each scenario, the best hyperparameters might differ. The interplay between hyperparameters and data traits is vital for model efficacy and adaptability.
@@
In the proposed method, we train for $200$ epochs with a batch size of $1024$ and a learning rate of $0.001$. The hyperparameters are optimized by grid search in the number of layers in vector field $n_{l}$ of $\left\{16,32,64,128\right\}$ and hidden vector dimensions $n_{h}$ of $\left\{16,32,64,128\right\}$. The best hyperparameters are highlighted with bold font in Table 15. Relative to the naïve Neural SDE, our suggested approaches deliver enhanced results through a more intricate model framework.
@@
We utilized the Python library torchcde [^12] https://github.com/patrick-kidger/torchcde (Kidger et al., 2020; Morrill et al., 2021b) for the interpolation scheme and the Python library torchsde [^13] https://github.com/google-research/torchsde (Li et al., 2020) to solve Neural SDEs for both the Neural SDE and the proposed methods. This study utilized the original Neural CDE source code [^14] https://github.com/patrick-kidger/NeuralCDE (Kidger et al., 2020) for benchmark methods including $\Delta t, GRU$-D, GRU-ODE, ODE-RNN, and Neural CDE. Additionally, we employed ODE-LSTM using its original source code [^15] https://github.com/mlech26l/ode-lstms (Lechner & Hasani, 2020), as well as Neural RDE [^16] https://github.com/jambo6/neuralRDEs (Morrill et al., 2021a), ANCDE [^17] https://github.com/sheoyon-jhin/ANCDE (Jhin et al., 2023b), EXIT [^18] https://github.com/sheoyon-jhin/EXIT (Jhin et al., 2022), and LEAP [^19] https://github.com/alflsowl12/LEAP (Jhin et al., 2023a), using their original source code. For Latent SDE, we formulate adjoint SDE proposed by Li et al. (2020) and train backpropagation incorporated with the original KL divergence.
@@
In each model and dataset, the following hyperparameters are optimized to minimize validation loss: hidden vector dimensions $n_{h}$, and the number of layers $n_{l}$. For RNN-based methods such as RNN, LSTM, and GRU, the fully-connected layer employs hyperparameters. For NDE-based methods, they are utilized to embed layer and vector fields. The hyperparameters are optimized as follows: the learning rate $lr$ from $10^{-4}$ to $10^{-1}$ using log-uniform search; $n_{l}$ from $\left\{1,2,3,4\right\}$ using grid search; and $n_{h}$ from $\left\{16,32,64,128\right\}$ using grid search. The batch size was selected from $\left\{16,32,64,128\right\}$ with respect to total data size. All models were trained for 100 epochs, and the best model was selected based on the lowest validation loss, ensuring that the chosen model generalizes well to unseen data. These design choices and training strategies aim to create a robust, well-performing model that effectively addresses the given classification tasks.

### v3

@@ 2 Related Work & Preliminaries
Let $(\Omega,{\mathcal{F}},P)$ be a probability space. Fix integer $d\geq 1$. For an ${\mathbb{R}}^{d}$-valued random variable $X$, its law is denoted by ${\mathcal{L}}(X)$. ||$\cdot$|| ⋅ | is the Euclidean norm where the dimension of the space may vary depending on the context. We denote by $L^{p}(\Omega)$ the set of $X$ with $(\mathbb{E}[|X|^{p}])^{1/p}<\infty$. Let ${\mathcal{P}}({\mathbb{R}}^{d})$ denote the set of probability measures on ${\mathcal{B}}({\mathbb{R}}^{d})$. For two probability measures $\mu,\nu\in{\mathcal{P}}({\mathbb{R}}^{d})$, let ${\mathcal{C}}(\mu,\nu)$ denote the set of probability measures $\Pi$ on ${\mathcal{B}}({\mathbb{R}}^{2d})$ such that its respective marginals are $\mu$ and $\nu$. Then, the Wasserstein distance of order $p\geq 1$ is defined as
@@ 3.1 The proposed Neural SDEs
where the initial condition $h(\cdot;\theta_{h})$ is an affine function with parameter $\theta_{h}$, the drift function $\gamma(\cdot,\cdot;\theta_{\gamma})$ is a neural network with parameter $\theta_{\gamma}$ and the diffusion function $\sigma(\cdot;\theta_{\sigma})$ is a neural network with parameter $\theta_{\sigma}$. Due to its exponential form, Neural GSDE has distinct characteristics associated with deep ReLU networks such as a unique positive solution and an absorbing state of 00, which are not observed in Neural LSDE and Neural LNSDE (see Proposition 3.4 (ii)).
@@ 3.2 Properties of the proposed Neural SDEs
    Assume that the activation function in $\gamma$ is either a tanh or sigmoid function. Then, the SDE in Equation (6) has a unique strong solution ${\bm{z}}(t)$, which is nonnegative almost surely, i.e., $P({\bm{z}}(t)\geq 0\mbox{ for all $t\geq 0$})=1$ a.s. Furthermore, state 00 is an absorbing state.
The proof for Proposition 3.4 can be found in Appendix A. Intuitively, Proposition 3.4 (ii) implies that the latent representation ${\bm{z}}(t)$ of the Neural GSDE has always nonnegative values. Furthermore, once the one of values of the latent representation reaches 00, it remains there forever. This property is consistent with that of ReLU networks which have positive values for activated neurons and zeros for deactivated neurons. In particular, the fact that deactivated neurons never turn to activated neurons corresponds to the property of state 00 being an absorbing state.
@@ 3.3 Robustness under distribution shift
Let ${\bm{x}}$ denote the input data, and its law be represented by ${\mathcal{L}}({\bm{x}})$, i.e., ${\bm{x}}$ is a ${\mathbb{R}}^{d_{x}}$-valued random variable. Consider $\widetilde{\bm{x}}$ to be a perturbed version of the input data, with the law ${\mathcal{L}}(\widetilde{\bm{x}})$, such that
@@
${\mathcal{W}}_{1}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}))$ $\leq\sqrt{3}L_{F}L_{h}c_{1}e^{-c_{2}T}\sqrt{\left(5+2\mathbb{E}[|{\bm{x}}|^{4}]+2\mathbb{E}[|\widetilde{\bm{x}}|^{4}]]\right)}\rho,$
${\mathcal{W}}_{2}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}))$ $\leq\sqrt{2\sqrt{3}L_{h}c_{1}}L_{F}e^{-c_{2}T/2}\left(5+2\mathbb{E}[|{\bm{x}}|^{4}]+2\mathbb{E}[|\widetilde{\bm{x}}|^{4}]]\right)^{1/4}\sqrt{\rho},$
@@ Results.
Methods Mean Squared Error (×$\times 10^{-3}$)
@@ Acknowledgement
where $\gamma_{i}$ is the $i$-th element of $\gamma$ and $(\sigma(t;\theta_{\sigma}))_{ij}$ represents the $(i,j)$-th element of $\sigma(t;\theta_{\sigma})$, and $[{\bm{y}}_{i}(t),{\bm{y}}_{i}(t)]$ is the quadratic variation of ${\bm{y}}_{i}(t)$ up to $t$ for $i=1,2,\ldots,d_{z}$. In the matrix form, ${\bm{z}}(t)$ satisfies
@@
Therefore, ${\bm{z}}(t)$ is the solution of Equation (6). Let $T_{0}^{i}=\inf\{t>0|z_{i}(t)=0\}$ be the first hitting time to state 00 with $z_{i}(0)>0$ for $i=1,2,\ldots,d_{z}$ where $z_{i}(t)$ is the $i$-th component of ${\bm{z}}(t)$. Since ${\mathrm{d}}z_{i}(t)=0$ for $t\geq T_{0}^{i}$, then $z_{i}(t)=0$ for all $t\geq T_{0}^{i}$. This implies that once $z_{i}(t)$ reaches 00, it remains there forever. Therefore, 00 is an absorbing state. ∎
@@
[…]
@@ B.1 Background on stability of SDEs
${\mathcal{W}}_{1}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}))$ $={\mathcal{W}}_{1}\left({\mathcal{L}}(F({\bm{z}}(T))),{\mathcal{L}}(F(\widetilde{\bm{z}}(T))\right)$
$=\inf_{\Pi\in{\mathcal{C}}}\left(\int_{{\mathbb{R}}^{d}}\int_{{\mathbb{R}}^{d}}|F({\bm{z}})-F(\widetilde{\bm{z}})|\Pi({\mathrm{d}}{\bm{z}},{\mathrm{d}}\widetilde{\bm{z}})\right)$
@@
$\leq L_{h}c_{1}e^{-c_{2}T}w_{1,2}({\mathcal{L}}({\bm{x}}),{\mathcal{L}}(\widetilde{\bm{x}}))$
$\leq L_{h}c_{1}e^{-c_{2}T}\mathbb{E}\left[(1\wedge|{\bm{x}}-\widetilde{\bm{x}}|)\left(1+V_{2}({\bm{x}})+V_{2}(\widetilde{\bm{x}})\right)\right]$
@@
Now focus on deriving the upper bound for ${\mathcal{W}}_{2}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}))$. Observe that
${\mathcal{W}}_{2}^{2}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}))$ $={\mathcal{W}}_{2}^{2}\left({\mathcal{L}}(F({\bm{z}}(T))),{\mathcal{L}}(F(\widetilde{\bm{z}}(T))\right)$
$=\inf_{\Pi\in{\mathcal{C}}}\left(\int_{{\mathbb{R}}^{d}}\int_{{\mathbb{R}}^{d}}|F({\bm{z}})-F(\widetilde{\bm{z}})|^{2}\Pi({\mathrm{d}}{\bm{z}},{\mathrm{d}}\widetilde{\bm{z}})\right)$
@@
$\leq 2L_{h}c_{1}e^{-c_{2}T}w_{1,2}({\mathcal{L}}({\bm{x}}),{\mathcal{L}}(\widetilde{\bm{x}}))$
$\leq 2L_{h}c_{1}e^{-c_{2}T}\mathbb{E}\left[(1\wedge|{\bm{x}}-\widetilde{\bm{x}}|)\left(1+V_{2}({\bm{x}})+V_{2}(\widetilde{\bm{x}})\right)\right]$
@@ B.3 Linear Noise SDE and Geometric SDE
${\mathcal{W}}_{1}({\mathcal{L}}({\bm{y}}),{\mathcal{L}}(\widetilde{\bm{y}}))$ $={\mathcal{W}}_{1}\left({\mathcal{L}}(F({\bm{z}}(T))),{\mathcal{L}}(F(\widetilde{\bm{z}}(T))\right)$
$=\inf_{\Pi\in{\mathcal{C}}}\left(\int_{{\mathbb{R}}^{d}}\int_{{\mathbb{R}}^{d}}|F({\bm{z}})-F(\widetilde{\bm{z}})|\Pi({\mathrm{d}}{\bm{z}},{\mathrm{d}}\widetilde{\bm{z}})\right)$ (24)
@@ D.2 Network Architecture
In practice, we implemented several design choices to optimize the model’s performance and stability. Firstly, we utilized the idea of sinusoidal positional encoding, as proposed by Vaswani et al. (2017), for the time variable, ensuring that each time step possesses a unique encoding. Secondly, we opted for the tanh nonlinearity as the final operation for drift, diffusion, and all other vector fields, following the recommendation by Kidger et al. (2020). This choice is intended to prevent the model from experiencing issues related to excessively large values or gradients. Lastly, we employed layer-wise learning rates for the model’s final layer (e.g. ×$\times 100$), facilitating more precise and adaptive learning for the classification task. This approach can result in improved classification outcomes, as each layer can adjust its learning rate according to the complexity of the features it aims to capture.
@@ D.3 Training strategy
5:     Train parameters $\bm{\theta}=\left[\theta_{\zeta},\theta_{\gamma},\theta_{\sigma},\theta_{\textit{MLP}}\right]$ using $D_{train}$ and classification loss $\mathcal{L}(y,\hat{y})$.
$\operatorname*{arg\,min}_{\bm{\theta}}\mathcal{L}(y,\hat{y}).$
@@ Appendix E Detailed results of the benchmark datasets
For the proposed methodology, the training process spans $300$ epochs, employing a batch size of $64$ and a learning rate of $0.001$. To train our models on a dataset comprising irregularly sampled time series, we adopt a strategy from Shukla & Marlin (2021). This involves the modified VAE training method, where we optimize a normalized variational lower bound of the log marginal likelihood, grounded on the evidence lower bound (ELBO). Hyperparameter optimization is conducted through a grid search, focusing on the number of layers in vector field $n_{l}$ of $\left\{1,2,3,4\right\}$ and hidden vector dimensions $n_{h}$ of $\left\{16,32,64,128\right\}$. The optimal hyperparameters for the proposed methods are remarked in bold within Table 10, encompassing observed data ranging from 50% to 90%.
@@
For the proposed methodology, we spanned $500$ epochs with a batch size of $1024$ and a learning rate set at $0.001$. We performed hyperparameter optimization using a grid search, focusing specifically on the number of layers in vector field $n_{l}$ of $\left\{1,2,3,4\right\}$ and hidden vector dimensions $n_{h}$ of $\left\{16,32,64,128\right\}$. The optimal hyperparameters, as identified for our methods, are highlighted in bold in Table 11 for regular datasets. For scenarios with 30%, 50%, and 70% missing data, we applied the same hyperparameters.
@@
In the proposed method, we train for $200$ epochs with a batch size of $1024$ and a learning rate of $0.001$. The hyperparameters are optimized by grid search in the number of layers in vector field $n_{l}$ of $\left\{1,2,3,4\right\}$ and hidden vector dimensions $n_{h}$ of $\left\{16,32,64,128\right\}$. The best hyperparameters are highlighted with bold font in Table 13 and 14. Due to varying input channel counts in each scenario, the best hyperparameters might differ. The interplay between hyperparameters and data traits is vital for model efficacy and adaptability.
@@
In the proposed method, we train for $200$ epochs with a batch size of $1024$ and a learning rate of $0.001$. The hyperparameters are optimized by grid search in the number of layers in vector field $n_{l}$ of $\left\{1,2,3,4\right\}$ and hidden vector dimensions $n_{h}$ of $\left\{16,32,64,128\right\}$. The best hyperparameters are highlighted with bold font in Table 15. Relative to the naïve Neural SDE, our suggested approaches deliver enhanced results through a more intricate model framework.
@@
We utilized the Python library torchcde [^12] https://github.com/patrick-kidger/torchcde (Kidger et al., 2020; Morrill et al., 2021b) for the interpolation scheme and the Python library torchsde [^13] https://github.com/google-research/torchsde (Li et al., 2020) to solve Neural SDEs for both the Neural SDE and the proposed methods. This study utilized the original Neural CDE source code [^14] https://github.com/patrick-kidger/NeuralCDE (Kidger et al., 2020) for benchmark methods including GRU-$\Delta t$, GRU-D, GRU-ODE, ODE-RNN, and Neural CDE. Additionally, we employed ODE-LSTM using its original source code [^15] https://github.com/mlech26l/ode-lstms (Lechner & Hasani, 2020), as well as Neural RDE [^16] https://github.com/jambo6/neuralRDEs (Morrill et al., 2021a), ANCDE [^17] https://github.com/sheoyon-jhin/ANCDE (Jhin et al., 2023b), EXIT [^18] https://github.com/sheoyon-jhin/EXIT (Jhin et al., 2022), and LEAP [^19] https://github.com/alflsowl12/LEAP (Jhin et al., 2023a), using their original source code. For Latent SDE, we formulate adjoint SDE proposed by Li et al. (2020) and train backpropagation incorporated with the original KL divergence.
@@
In each model and dataset, the following hyperparameters are optimized to minimize validation loss: the number of layers $n_{l}$ and hidden vector dimensions $n_{h}$. For RNN-based methods such as RNN, LSTM, and GRU, the fully-connected layer employs hyperparameters. For NDE-based methods, they are utilized to embed layer and vector fields. The hyperparameters are optimized as follows: the learning rate $lr$ from $10^{-4}$ to $10^{-1}$ using log-uniform search; $n_{l}$ from $\left\{1,2,3,4\right\}$ using grid search; and $n_{h}$ from $\left\{16,32,64,128\right\}$ using grid search. The batch size was selected from $\left\{16,32,64,128\right\}$ with respect to total data size. All models were trained for 100 epochs, and the best model was selected based on the lowest validation loss, ensuring that the chosen model generalizes well to unseen data. These design choices and training strategies aim to create a robust, well-performing model that effectively addresses the given classification tasks.